OpenStudy (anonymous):

2/5 divided by 5 = ? 1/8 divided by 6= ? 4/6 divided by 2 = ? 9/5 divided by 3 = ? 10/5 divided by 4 = ? 5/2 divided by 2 = ?

OpenStudy (anonymous):

Medals for answers

OpenStudy (anonymous):

the first one is 0.08

OpenStudy (anonymous):

I need them as fractions.

OpenStudy (anonymous):

2/25

OpenStudy (anonymous):

2/25 for which one?

OpenStudy (anonymous):

1/48(second)

OpenStudy (anonymous):

4/6 divided by 2: 1/3

OpenStudy (anonymous):

9/5 divided by 3: 3/5

OpenStudy (anonymous):

10/5 divided by 4 : 1/2

OpenStudy (anonymous):

5/2 divided by 2 : 5/4

OpenStudy (anonymous):

2/25 the first one

OpenStudy (anonymous):

done:)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

2/5 divided by 7 = ?

OpenStudy (anonymous):

\[{2\over5}\times{1\over5} = {2\over25}\]\[{1\over8}\times{1\over6} = {1\over48}\]\[{4\over6}\times{1\over2} = {4\over12} = {1\over3}\]\[{9\over5}\times{1\over3} = {9\over15} = {3\over5}\]\[{10\over5}\times{1\over4} = {10\over20} = {1\over2}\]\[{5\over2}\times{1\over2} = {5\over1} = 5\]

OpenStudy (anonymous):

\[{2\over5}\times{1\over7} = {2\over35}\]